1.A Spherical body of 2.0 cm diameter is maintained at 600 0C. Assuming that it radiates as if it were a black body (perfect emitter)?

(i) Determine the surface area A, of the spherical body.

(ii) At what rate in watts is energy radiated from the sphere?

The radiant energy Q, emitted in a time t by an object that has a kelvin temperature T, a surface area A, and emissivity e, is given by

Q = euT^4 At and Q/t = euT^4 A
where the temperature is in Kelvin.
1
Expert's answer
2018-08-27T09:43:09-0400
the surface area of the spherical body
A=4πR^2=4×3.14×〖0.01〗^2=1.256×〖10〗^(-3) m^2
The Stefan–Boltzmann law states
Q/t =eσT^4 A=1×5.67×〖10〗^(-8)×(600+273)^4×1.256×〖10〗^(-3)=41.4 W